A ladder 26 ft long is placed 10 ft from the base of a house. How high up the side of the house does the ladder reach?
Answer by mananth(16949)

let the height be x
using pythagoras theorem
x^2+10^2=26^2
x^2+100=676
x^2=576
x= sqrt576
x=24 feet the height